OpenStudy (anonymous):

Medal and fan! 1. 6 - x - x = 18 A: -6 B: 6 C: 9 D: -9 2. 4x - 3/5 = 2/5 A. 1/5 B: 5/5 C: 1/2 D. 1/4

OpenStudy (anonymous):

the first one cannot be defined

OpenStudy (anonymous):

sure it can

OpenStudy (anonymous):

\[6-x-x=18\] rigtht?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

on planet earth most people write \(-x-x\) as (-2x\) so start with \[6-2x=18\]

OpenStudy (anonymous):

then subtract \(6\) from both sides after that, divide both sides by \(-2\)

OpenStudy (anonymous):

So it would be B ( 6) right?

OpenStudy (anonymous):

oh yeah, im sorry did't see the minus sign there

OpenStudy (anonymous):

What about the second question? @satellite73

OpenStudy (dayakar):

4x - 3/5 = 2/5 4x-3/5+3/5 = 2/5+3/5 (adding both sides 3/5) 4x= 2/5+3/5 4x = 5/5 4x= 1 x= 1/4

OpenStudy (dayakar):

D is the answer

OpenStudy (anonymous):

Thank You.

OpenStudy (dayakar):

your welcome